TigerSharkie:

Which of the following correctly uses the absolute value to show the distance between −70 and 15?

russianmafiya:

absolute value means the apposite so like 23 its absolute value would be -23

TigerSharkie:

answers |−70 − 15| = |−85| = −85 units |−70 + 15| = |−55| = 55 units |−70 − 15| = |−85| = 85 units |−70 + 15| = |−55| = −55 units

TigerSharkie:

wait so which one is it

russianmafiya:

so think about it 15 is positive so an other way to write it is +15

TigerSharkie:

so its either number 2 or 4??

russianmafiya:

okay yes so you slashed the trash now what is -70 + 15

TigerSharkie:

-55?

russianmafiya:

okay so now both of them are -55 but one of them is correct so now what is the absolute value of -55

TigerSharkie:

-55

russianmafiya:

basalt value is the opposite of the number

russianmafiya:

like the absolute value of -99 is 99

TigerSharkie:

oh, so it would just be 55

russianmafiya:

yes so it would be the second one did you get it ?

TigerSharkie:

It doesn't say it's correct till the end of the lesson and no its not a test :p

TigerSharkie:

Im gonna close this one k cause Ima post my other one

russianmafiya:

well as long as you know how to do it your good

TigerSharkie:

Thank you!

russianmafiya:

your welcome